A 70 kg person does a bungee jump from a bridge. The natural length of the bungee cord is 20m and it has a spring constant k of 80 N/m. When the bungee cord has a total length of 25m, what is the acceleration of the jumper (to 2 significant figures)?
Answer #1

Use Newton’s law of motion to find the equation of motion for the person and solve that to find the acceleration as shown belowextension A2 = 25m - 20m - 5m - Fs - ung = ma - KAx-mg = ma => (80N/m) (5m) – (708) 9 = (709) a > a = - 4.1 m/s2 So, its 4.1

***********************************************************************************************
Check the answer and let me know immediately if you find something wrong or missing... I will rectify the mistakes asap if any